# Visual and auditory insights into multifrequency variability

### v2

@@ 3.2 Production
Once the processed data were available, the production consisted of converting signal to sound. For this, the data was normalised to be expressed in linear units with Musical Instrument Digital Interface (MIDI) notes.
Sonification was achieved using the Parameter Mapping Sonification (PMSon) with the technique ‘one-to-one’ mapping (Grond & Berger, 2011), where the parameters are mapped to the signal domain at a tempo of 80 bpm, C3 octave, using a relative d minor scale.
All the programming was carried out using the MIDItime 1.1.3 (Mark, 2016) package using the Python 3 programming language.
@@ 3.3 Post-production
The PMson used was the technique ‘one-to-many’ mapping or divergent mapping, which converts the MIDI sound to tinkle bell. This method expands the scales in the ranges of the perceptual domain.
The multifrequency sonificated objects present their time series, their waveforms and their spectrograms. Their auditory result can be consulted in the following website: Datasonification Project.
@@ 3.3.1 Waveforms and spectrograms
The visual representation of the objects were represented in three ways: the light curves, the waveforms and the spectrograms. All plots are paired in the dimension of time.
For the data of each blazar, these waveforms, are a representation of a sound wave in function of time, where the amplitude represents the volume of the sound. For all studied blazars, since the datasets were normalised, then the maximum amplitude is 1.
The spectrograms provide information about the frequencies that make up a sound, in which we can appreciate the graphic representation of how the frequency varies over time. The vertical color bar to the right of the spectrogram indicates the density accumulation of points given by the signal at each point in time and frequency. In this context, density accumulation refers to how strong the signal is—either in terms of photon counts or flux—at a given energy (or frequency) and time. Warmer colors (e.g., red or yellow) represent higher densities, meaning stronger emission, while cooler colors (e.g., blue or black) indicate lower densities or weaker signals. This allows us to visualize how the energy output of the source evolves over time and across different frequencies. Understanding the frequency, as the pitch of the instrument (in this case a tinkle bell instrument), the faster the wave repeats itself (given the frequency), the higher the pitch of the note.
It is important to mention that any physical object or phenomenon that emits noise periodically (hundreds or thousands of times per second) must have a perceptible tone. However, not all sound is periodic. For example, since white noise represents static sound, it is just a uniform distribution of audible frequencies, or Gaussian distribution. Therefore, because white noise is non-periodic, it does not have a noticeable pitch.
@@ 4 Discussion
Using a bell as a sonification instrument in this part of the production enhances clarity and auditory distinction. Its resonant tone makes it easier to perceive important events in the data without blending with other sounds.
Additionally, its sharp attack (well-defined onset) allows for precise marking of specific points in time, which is particularly useful for sonifying time series or discrete events. The bell also generates harmonics that enrich the tonal representation of data variations, helping to distinguish between different values or categories.
For example, in the light curves presented here, higher flux values translate into a more intense or higher-amplitude sound that gradually fades over time, fully covering the gaps in the curve. While these changes are evident in waveforms, they may go unnoticed in light curves. However, the ear can detect abrupt shifts in intensity or volume.
This is the advantage of sonification: what escapes the eye can be perceived by the ear.
@@
The waveform, on the other hand, captures the full temporal evolution of the signal with high amplitude resolution. It acts as a detailed zoom into the temporal variations that might not be visible in the light curve. This allows for the analysis of precise dynamics, such as rapid pulses, subtle oscillations, or specific shapes. It also helps to distinguish noise from real signals and correlate light curve peaks with their exact amplitude and timing.
The spectrogram decomposes the signal into its frequency components, showing how they evolve over time. It helps to identify events linked to specific frequency bands, such as peaks or oscillation modes and distinguishes noise, characterised by specific spectral patterns, from significant signals.
@@
  • The light curve offers a general view of the signal’s behaviour over time.
  • The waveform adds fine detail about amplitude variations.
  • The spectrogram reveals frequency properties, uncovering patterns not evident in the other two.
This integrated approach enables the correlation of complex events. For instance, a peak in the light curve can be examined in the waveform to determine if it represents a sudden event and in the spectrogram to check for associated frequency changes. This is relevant for identifying specific physical phenomena, such as periodic pulses, spectral bursts, or emission changes, often observed in the study of blazars or binary stars.
@@
For example, in the observation of a blazar, the light curve tracks global brightness variations, the waveform reveals rapid underlying pulses invisible on a larger scale, and the spectrogram identifies harmonic components or specific spectral features.
The synchronised combination of light curve, waveform, and spectrogram provides a comprehensive and integrated view of the data: intensity, temporal dynamics, and frequency structure. This enhanced analysis, facilitates the identification of complex events and improves the interpretation of physical phenomena.
As mentioned in the Introduction, there are already several works on data sonification in astronomy. These works provide important information from the perspective of sonifying the image, since the areas with the highest light intensity can be distinguished with the ear. In addition, they have a high pedagogical value, since they allow the public to have a didactic and playful approach to astronomy.
However, these works do not show on the time line how are the variations, trends, cyclicity, periodicity, seasonality of the phenomena or of the radiative processes. In this sense, the route that we have taken is that of sonification of the time series, or light curves, and the exercise that we have carried out has been on nine active galaxies that contain a supermassive black hole in their nucleus, an accretion disk that surrounds it, and a relativistic jet, so the radiation emission from these objects is multifrequency. Thus, sonification data for the Mrk 501, Mrk 1501, Mrk 421, BL Lacerta, AO 0235+164, 3C 66A (PKS 0219+428), OJ 049 (PKS 0829+046), OJ 287, and PKS J2134-0153 blazars were presented with their associated light curves, waveforms and spectrograms.
@@ 5 Conclusions
Although this approach does not aim to definitively demonstrate periodicities, the auditory representation offers an innovative way to identify potential patterns, correlate events across different time scales, and explore the complex dynamics of these astrophysical objects. This work lays the groundwork for future research to delve deeper into the potential of sonification to highlight regularities and transient phenomena in astronomical data.
This multimodal approach contributes to a more inclusive and diverse understanding of blazar behaviour, paving the way to expand scientific analysis through methods that integrate both visual and auditory perception.
The radiation processes in the studied galaxies are multifrequency, so the data we have processed for sonification correspond to radio, optical, X-ray, and gamma-ray bands. Thus, the auditor can listen to the recorded signals that are emitted due to the radiative processes that distinguish the components of the active nucleus of the galaxy. Data were obtained from seven observatories: UMRAS, AAVSO, Astrogeo-VLBI, Swift, Fermi, and ZTF.
The graphs presented for each of the blazars are specially useful for detecting changes over time, as well as getting a closer idea of sonification in visual terms. In these one can visualise what has been sounded, the changes in intensity (volume) and frequency (tone).
It is important to say that other fields in astronomy in which the sonification of data is present are in matters of education and public participation. There is a record of how sonifying images is impacting the dissemination of astronomy, and how citizen participation is increasing (Zanella et al., 2022).
The sonification of data is a way with which substantial information can be extracted, such as the search for periodicities, power increases, regularities, absences of information, etc. In addition to visualising the data, we have sonified it with the purpose of disseminating our advances to an even broader spectrum of society, at the same time that we investigate auditory events through the sense of hearing that help to distinguish periodicities.
Thus, the central interest of this work has been to disseminate the analysis of astrophysical data from a technique that is attractive to science in general, which is listening to the phenomena. In this sense, this is a work to bring the general public closer to astrophysics, from the sonification of the light curves of the active nuclei of galaxies, but also from the concern to socially include in science sectors of the population are blind or vision impaired (BVI) (Pérez-Montero, 2019).

### v3

@@ 3.2 Production
Once the processed data were available, the production consisted of converting signal to sound.
For this purpose, each dataset was first normalized using a linear min-max scaling:
$\textrm{normalized value}=\frac{\text{flux}-\min(\text{flux})}{\max(\text{flux})-\min(\text{flux})},$ (1)
where flux is the measured brightness at a given time, and $\min(\text{flux})$ and $\max(\text{flux})$ correspond to the minimum and maximum flux values in the dataset. This procedure ensures that all data points are rescaled to the $[0,1]$ interval, facilitating their subsequent mapping into musical notes.
The normalized values were then mapped to Musical Instrument Digital Interface (MIDI) notes using Parameter Mapping Sonification (PMSon) with the ‘one-to-many’ mapping technique (Grond & Berger, 2011). In this approach, each normalized data point triggers a primary note (encoding flux as pitch) and a set of characteristic overtones, producing a richer auditory representation. Additionally, the amplitude (loudness) of each note is directly tied to the flux, meaning that the same physical quantity is represented both in pitch and amplitude. This dual encoding enhances the perception of variations, trends, and transient events in the light curves.
To map the normalized values to the notes of the scale, the interval [0, 1] was divided into seven equal bins, each corresponding uniquely to one of the seven notes of the D minor scale within a single octave. For example, a normalized value of 0.0 corresponds to the first note (D), a value of 0.5 corresponds to the fourth note (G), and a value of 1.0 to the highest note (C) in the octave.
The sonification was implemented in the C3 octave, with a tempo of 80 beats per minute (bpm), where each note represents one time step of the light curve. In this context, the 80 bpm provides a temporal spacing between successive notes, corresponding to the time evolution of the observed flux. Different octave ranges were assigned to each wavelength band during post-production. This choice does not imply any physical relationship between electromagnetic wavelength and pitch; rather, it was a perceptual design decision intended to enhance audibility and facilitate differentiation between bands in the multiwavelength sonification.
Because the data span different time intervals and have varying sampling rates across wavelength bands, the light curves were rescaled in time for the sonification. Successive data points were mapped to fixed note durations at 80 bpm, representing the relative temporal evolution of each dataset. Absolute timing (MJD) is not preserved in the sonification; instead, temporal structures and patterns within each wavelength are maintained. This allows auditory perception of variability in each band, while the differences in time span and sampling between bands provide complementary insights. Key intervals, including known quasi-periodicities, are annotated in the figures and on the project web page to guide the listener in interpreting the sonified data.
The programming was carried out using the MIDItime 1.1.3 package (Mark, 2016) in Python 3. Data uncertainties were not directly encoded in the sonification, although they informed the preprocessing stage by guiding the removal or correction of anomalous points prior to mapping.
@@ 3.3 Post-production
The tinkle bell produces a percussive, bell-like sound with a fundamental pitch and a set of inharmonic overtones (i.e., partials that are not exact integer multiples of the fundamental frequency). Each note has a rapid attack and an exponential decay determined by the instrument’s envelope. The spectral components scale proportionally with the input pitch, so higher notes produce perceptually brighter sounds. This instrument was chosen because its clear and transient timbre and distinctive spectral structure allow individual data points to be audibly distinguished, even in dense sequences.
Parameter Mapping Sonification (PMSon) was applied using the ‘one-to-many’ mapping technique. In this approach, a single normalized data point generates multiple simultaneous partials based on the tinkle bell’s harmonic structure. Specifically, each normalized data value is first mapped to a primary MIDI note; this note then triggers the instrument’s characteristic overtones, producing a richer auditory representation and expanding the perceptual pitch content. To ensure perceptual separation between wavelength regimes (radio, optical, X-ray, $\gamma$-ray), each band was assigned to a distinct octave range via pitch offsets applied after the mapping stage. This strategy preserves the relative variations within each light curve while allowing the different regimes to occupy clearly distinguishable pitch regions.
Each note has a constant duration of one beat at 80 bpm, corresponding to the temporal spacing defined in the sonification. The tinkle bell instrument was configured with an exponential decay of approximately 0.5 seconds, allowing individual notes to be clearly perceived while avoiding excessive overlap. Temporal overlap naturally occurs if data points are closer than the beat interval. Gaps in the light curves are represented by silent intervals of corresponding length, preserving the temporal structure of the original data.
The multifrequency sonified objects are illustrated in figures showing:
  1. 1.
    the original light curve (time series),
  2. 2.
    the sonified waveform, and
  3. 3.
    the corresponding spectrogram, allowing visualization of both temporal and frequency content.
The auditory results of these sonifications can be accessed at the Datasonification Project website.
@@ 3.3.1 Waveforms and spectrograms
The visual representation of the objects were represented in three ways: the original light curves, the sonification waveforms and the spectrograms. All plots are aligned along the time axis for direct comparison.
For each blazar, the waveforms represent the audio signal as amplitude versus time, where amplitude corresponds to the perceived loudness of each note. Since the datasets were normalized prior to sonification, the maximum amplitude in all waveforms is 1. Importantly, the waveforms do not encode pitch; they only show how the loudness of the sound varies over time.
The spectrograms were generated from the sonified audio rather than from the original light curve data. They display the distribution of frequencies in the sonification as a function of time. The vertical color bar indicates the intensity of the signal in each time–frequency bin. Warmer colors (e.g., red or yellow) correspond to stronger signal components, while cooler colors (e.g., blue or black) indicate weaker ones. The presence of multiple bright frequencies within a single time bin arises both from the harmonic overtones characteristic of the tinkle bell instrument and from temporal overlap of successive notes.
Together, the light curves, waveforms, and spectrograms provide complementary perspectives on the same data: the original flux variations, the time-domain structure of the sonification, and the distribution of its pitch components across time.
@@ 4 Discussion
Using a bell as a sonification instrument in this part of the production enhances clarity and auditory distinction. Its resonant tone, rich in overtones, has been shown to facilitate the auditory salience of events by making them stand out from the background (Kramer et al., 1999; Grond & Berger, 2011). Previous studies in auditory display and sonification highlight that instruments with clear attack and sustained resonance improve the perceptual separation of events, reducing masking effects and allowing listeners to identify patterns more easily (Hermann, 2008). Therefore, the tinkle bell timbre was chosen to ensure that important variations in the data could be perceived without blending with other sounds.
Additionally, its sharp attack (well-defined onset) allows for precise marking of specific points in time, which is particularly useful for sonifying time series or discrete events. The bell timbre is also characterized by a rich harmonic spectrum: when a note is struck, besides the fundamental frequency, several overtones of different amplitudes are produced. These overtones increase the perceptual brightness and richness of the sound, which improves the listener’s ability to discriminate subtle variations between consecutive notes (Zwicker & Fastl, 1999). In practice, this means that small changes in the mapped data values are more easily perceived as distinct auditory events, as the harmonics reinforce pitch perception and provide additional auditory cues beyond the fundamental tone.
For example, in the light curves presented here, the amplitude of each note is directly tied to the flux value of the corresponding data point: higher flux corresponds to a higher amplitude. The note length is constant for all points and is chosen to ensure that each note sufficiently covers the temporal spacing between consecutive measurements, effectively bridging gaps in the light curve. This approach allows abrupt changes in intensity or volume to be perceived by ear, even if they are not visually evident in the waveform or the original light curve. Consequently, sudden increases or decreases in flux are more perceptible auditorily, providing an additional modality for identifying significant variations in the data.
This is the advantage of sonification: features that may be difficult to perceive visually, such as rapid short-term flares or subtle oscillations in the flux, can be detected by the ear. When sonified, these variations correspond to sudden changes in pitch or amplitude, allowing listeners to perceive patterns and events that might escape visual inspection of the light curves.
@@
The waveform captures the full temporal evolution of the signal, providing a zoom into temporal variations that might not be visible in the light curve. For example, the oscillations in Mrk 501 that appear as small irregularities in the light curve allow us to separate short-term variations from instrumental or observational noise. Similarly, the sonification and waveform visualization make it possible to correlate the peaks in the light curve with their amplitude and timing, such as the 224-day periodicity observed in Mrk 501 (Magallanes-Guijón, 2020), highlighting both the intensity and the temporal spacing.
The spectrogram is computed from the sonified audio signal rather than from the original astronomical data. It decomposes this audio waveform into its frequency components, showing how the pitch content generated by the sonification evolves over time. For example, in the case of PKS J2134-0153 (Figure 7), short-lived X-ray and $\gamma$-ray flares that appear only as subtle bumps in the light curve emerge as transient bright bands in the spectrogram. Because the tinkle bell instrument produces a fundamental pitch together with inharmonic overtones, each mapped data event spreads across multiple frequency bands, which become visible in the spectrogram. This visualization therefore reveals the temporal distribution, harmonic structure, and overlap of sonified events, complementing both the light curve and the auditory perception.
@@
  • The light curve offers a general view of the signal’s behaviour over time. For example, in OJ 287, the quasi-periodic variability is visible is visible in the light curve as a series of recurrent peaks, establishing the large-scale context of variability.
  • The waveform adds fine detail about amplitude variations. Rapid flares that appear only as small bumps in the light curve become clearly distinguishable in the waveform, allowing us to separate genuine short-term variability from instrumental noise.
  • The spectrogram reveals frequency properties, uncovering patterns not evident in the other two. In the case of Mrk 421, the use of the tinkle bell instrument generates harmonics that appear as multiple bright bands in the spectrogram. While this complicates the auditory perception, it enhances the visual identification of event timing and harmonic structure, making subtle features—such as overlapping flares—more apparent.
This integrated approach enables the correlation of complex events. For instance, a peak in the light curve can be examined in the waveform to determine if it represents a sudden event and in the spectrogram to check for associated frequency changes. This is relevant for identifying specific physical phenomena, such as periodic pulses, spectral bursts, or emission changes, often observed in the study of blazars or binary stars. The corresponding spectrogram of Mrk 501, for example, shows transient bright frequency optical band, indicating the presence of oscillations.
@@
For instance, in case of OJ 049 a small fluctuation in the light curve initially appears as a possible flare. However, inspection of the waveform shows no coherent pulse dynamics, and the spectrogram lacks any associated frequency bands, confirming that the fluctuation is consistent with observational noise. The light curve reveals the overall flare duration and peak intensity, the waveform captures the sharp rise and decay dynamics, and the spectrogram highlights transient frequency components produced by the sonified signal.
This integrated analysis of OJ 049 not only confirms which apparent flares are artifacts of noise but also reveals genuine variability signatures. In particular, during epochs of increased activity, the waveform displays brief amplitude surges that correspond to microflares lasting only a few hours, while the spectrogram shows simultaneous enhancements across multiple frequency bands. These correlated features indicate that the variability arises from real physical changes-most likely turbulent processes or shock propagation within the jet-rather than from instrumental or sampling effects. Thus, combining the three representations enables a clearer discrimination between stochastic noise and intrinsic variability in the blazar’s emission.
The synchronised combination of light curve, waveform, and spectrogram provides a comprehensive and integrated view of the data: intensity, temporal dynamics, and frequency structure. This enhanced analysis, facilitates the identification of complex events and improves the interpretation of physical phenomena. In the case of OJ 049, the joint examination of the three representations reveals that apparent fluctuations in the optical light curve without corresponding features in the waveform or spectrogram are consistent with observational noise. Conversely, when a true flare occurs, the waveform shows coherent amplitude rises and decays, and the spectrogram displays transient frequency bands linked to the event. Similarly, in Mrk 421, correlated peaks across both optical and X-ray spectrograms highlight multiwavelength variability episodes, confirming that the combined visual–auditory approach enhances the detection and understanding of real astrophysical processes.
As mentioned in the Introduction, there are already several works on data sonification in astronomy, which explore different mapping strategies and applications. For instance, Dubus & Bresin (2013) provided a systematic review of sonification mapping strategies for physical quantities, establishing a foundation for perceptual and analytical approaches. More recent studies, such as Guiotto Nai Fovino et al. (2024), evaluated the effectiveness of sonification for time-series data exploration, highlighting its potential to enhance data comprehension beyond visual analysis. Likewise, Tucker Brown et al. (2022b) tested the efficacy of sonification for signal detection in light curves through the ASTRONIFY framework, demonstrating measurable improvements in identifying transient phenomena.
From a broader perspective, Mis examined how sound experts perceive astronomy sonification projects, emphasizing both their scientific and communicative dimensions. Similarly, Casado & García (2024) discussed multimodal analysis approaches, integrating auditory and visual representations to strengthen the interpretative process of astrophysical data. These works not only confirm the analytical value of sonification but also underscore its pedagogical potential (Zanella et al., 2022), as it enables the public to experience complex astronomical data in an engaging manner.
While previous works on sonification have focused primarily on mapping static data or images, they generally do not display how variations, trends, cyclicity, periodicity, or seasonality evolve over time. In contrast, our approach focuses on the sonification of temporal data – specifically, light curves – to capture the dynamic behavior of astrophysical phenomena.
We applied this method to nine active galaxies hosting supermassive black holes with surrounding accretion disks and relativistic jets, whose radiation exhibits multifrequency variability. The sources analyzed include Mrk 501, Mrk 1501, Mrk 421, BL Lacertae, AO 0235+164, 3C 66A (PKS 0219+428), OJ 049 (PKS 0829+046), OJ 287, and PKS J2134-0153. For each case, we present the sonified data alongside their corresponding light curves, waveforms, and spectrograms, allowing for a multiscale and multimodal interpretation of their emission behavior.
@@ 5 Conclusions
In addition, these sonifications serve a dual purpose for the visually impaired community. First, they provide a means for BVI scientists to perceive patterns, periodicities, and transient events directly through auditory cues, supporting independent data exploration. Second, they can be used as an educational and communicative tool, enabling the broader BVI audience to access and understand complex astrophysical phenomena. For example, short-lived oscillations that appear as subtle fluctuations in a light curve can become more perceptually salient in the sonification, allowing listeners to detect features that are otherwise difficult to visualise.
@@
Although this study does not aim to definitively demonstrate periodicities, the auditory representation provides an approach to detect potential patterns, correlate events across different time scales, and explore the complex dynamics of these astrophysical objects. By demonstrating how sonification can complement traditional visualizations, this work establishes a methodological foundation for future research, including systematic searches for periodicities, transient events, and multiwavelength correlations in astronomical data.
This multimodal approach enables a more inclusive and diverse understanding of blazar behaviour by allowing data to be explored both visually and auditorily. For example, visually impaired (BVI) scientists can access the temporal dynamics, amplitude variations, and frequency patterns in light curves through sonification, while sighted researchers can combine waveforms and spectrograms with sound to identify subtle features that might be missed visually. This integration enhances accessibility and provides complementary perspectives on the same data, improving the detection of flares, oscillations, and other transient phenomena.
The radiation processes in the studied blazars are multi-frequency, including synchrotron emission, inverse Compton scattering, and thermal radiation from the accretion disk. These processes produce variability in different energy bands (radio, optical, X-ray, and gamma-ray), which we sonified. As a result, the listener can perceive changes associated with specific radiative mechanisms: for example, rapid optical flares generated by synchrotron emission correspond to short, high-pitched notes, while slower variations in gamma-ray flux appear as longer, lower-pitched tones. In this way, the sonification provides an auditory representation of the physical processes driving the emission, linking observable features to their underlying astrophysical origins.
The graphs presented for each blazar are particularly useful for understanding the sonification in visual terms. By pairing the visualizations with the corresponding sounds, one can see how changes in flux or luminosity are reflected in the amplitude (volume) and frequency (pitch) of the sonified signal. This visual-auditory correspondence aids in interpreting complex temporal patterns, verifying features detected by ear, and providing a more comprehensive understanding of the data.
The sonification of data allows the extraction of substantial information, such as searching for periodicities, detecting increases in power, identifying regularities, or noting absences of data. Beyond visualizing the data, we sonified it to extend the accessibility of our results to a broader audience. In particular, this includes both the general public, who can gain an engaging and intuitive understanding of astrophysical phenomena through sound, and blind or visually impaired (BVI) (Pérez-Montero, 2019). individuals, including both laypeople and professional astronomers. By translating complex temporal patterns into auditory cues, the sonifications provide a pathway for BVI audiences to perceive and interpret features that might otherwise rely solely on visual inspection.
@@ Acknowledgements
Appendix A Blazar Sample and Specifications
For each dataset, the time resolution and integration per data point are determined by the corresponding instrument or database. In the optical band, ZTF data consist of single exposures of 30 s (Masci et al., 2019), while AAVSO records typically correspond to integrations of 60–300 s depending on the observer (Kafka & Beck, 2021). In the radio band, UMRAO data points were obtained from integrations of approximately 30 minutes per observation (Aller et al., 1985), and Astrogeo VLBI measurements correspond to individual interferometric sessions (Petrov et al., 2021). For the X-ray band, Swift-XRT provides observations with effective exposures ranging from a few hundred to a few thousand seconds (Burrows et al., 2005). In the case of the $\gamma$-ray band, Fermi-LAT light curves are generally binned into one-week intervals to ensure sufficient photon statistics (Ballet et al., 2023). These values represent the effective time resolution of the light curves used in this work.
Each blazar selected had the following specifications:
[…]
(7) PKS J2134-0153 has a redshift $z$ = 1.285, corresponding to a distance $\sim$ 5 Gpc (Truebenbach & Darling, 2017) with R. A. = 21h 34m 10.3095s and Dec. = $01^{\circ}$ 53’ 17.238” (Fomalont et al., 2000).
Appendix B Tables
In Tables 1 and 2, we use the following conventions: a Records refers to a single photometric or flux measurement in a given band, typically corresponding to one integration by the instrument or observer. The term Interim [d] indicates the total time span covered by the dataset, expressed as years, months, and days [y, m, d] between the first and last observation date, regardless of gaps. The column Total [d] refers to the number of distinct calendar days within that interval for which at least one measurement was obtained.
Blazar R.A. Dec. Redshift Distance
Mrk 501 16h53m52.2s +39°45’37” 0.034 140 Mpc
Mrk 1501 00h10m +10°58’ 0.089338 377 Mpc
BL Lacerta 22h02m43.3s +42°16’40” 0.07 0.276 Gpc
AO 0235+164 02h35m52.6s +16°24’05” 0.940 3 Gpc
3C 66A 02h23m12s +43°00’07” 0.34 1.4 Gpc
OJ 049 08h31m48.8s +04°29’39” 0.18 760 Mpc
PKS J2134-0153 21h34m10.3095s -01°53’17.238” 1.285 5 Gpc
Mrk 421 11h04m27.31s +38°12’31.8” 0.031 130 Mpc
OJ 287 08h54m48.87s +20°06’30.6” 0.306 1 Gpc
Table 1: Summary of the nine blazars analyzed in this work, including equatorial coordinates (R.A., Dec.), redshift, and approximate distance.
Blazar Band Records Interim Total Dates
[y, m, d] [d] [d]
Mrk 501 Optical 11,849 23, 2, 17 8,441 1998-06-24 – 2021-09-12
X-rays 28,000 12, 7, 9 4,607 2008-10-02 – 2021-10-08
$\gamma$-rays 4,199 11, 7, 9 4,239 2008-08-04 – 2020-03-13
Mrk 1501 Optical 1,255 11, 4, 0 4,141 2010-07-25 – 2021-11-25
X-rays 11,150 12, 5, 20 4,554 2008-10-02 – 2021-03-22
$\gamma$-rays 5,146 14, 2, 18 5,192 2008-08-04 – 2022-10-22
BL Lacertae Radio 413 12, 5, 25 4,560 2009-09-01 – 2012-05-21
Optical 10,117 52, 2, 27 19,082 1969-12-14 – 2022-03-13
X-rays 100,418 14, 1, 8 5,152 2008-10-02 – 2022-11-10
$\gamma$-rays 4,892 13, 7, 7 4,967 2008-08-04 – 2022-03-11
AO 0235+164 Optical 196 22, 9, 17 8,055 1998-08-25 – 2020-09-13
X-rays 61,146 9, 2, 5 3,353 2008-10-02 – 2017-12-07
$\gamma$-rays 4,223 11, 8, 1 4,262 2008-08-04 – 2020-04-05
3C 66A Radio 101 23, 4, 22 8,545 1997-01-11 – 2020-06-04
Optical 8,470 28, 0, 9 10,236 1994-02-01 – 2022-02-10
X-rays 3,163 14, 2, 7 5,181 2008-10-02 – 2022-12-09
$\gamma$-rays 4,223 11, 8, 1 4,262
OJ 049 Optical 76 17, 5, 22 6,383 2008-01-09 – 2020-06-23
X-rays 16,111 15, 2, 8 5,548 2002-10-09 – 2020-03-31
$\gamma$-rays 4,196 11, 8, 1 4,262 2008-08-05 – 2020-04-06
PKS J2134-0153 Optical 73 0, 10, 10 314 2009-10-04 – 2010-08-14
X-rays 4,123 6, 0, 6 2,197 2008-10-02 – 2014-10-08
$\gamma$-rays 4,223 11, 8, 1 4,262 2008-08-04 – 2022-10-22
Mrk 421 Optical 18,167 35, 9, 20 13,079 1984-06-22 – 2020-04-13
X-rays 160,346 15, 1, 0 5,510 2005-03-01 – 2020-04-01
$\gamma$-rays 4,129 11, 9, 3 4,294 2008-06-05 – 2020-03-08
OJ 287 Optical 13,754 32, 8, 6 11,939 1987-09-26 – 2020-06-03
X-rays 9,464 11, 5, 16 4,185 2008-11-12 – 2020-04-28
$\gamma$-rays 4,410 12, 2, 2 4,446 2008-04-04 – 2020-06-06
Table 2: Summary of all electromagnetic bands studied for the nine blazars. Records corresponds to a single flux measurement. Interim [d] is the total time span covered by the observations, expressed in years, months, and days [y, m, d]. Total [d] indicates the number of distinct days within the interim for which at least one measurement is available.